# Targeting Alignment: Extracting Safety Classifiers of Aligned LLMs

### v1

@@ Abstract
Alignment in large language models (LLMs) is used to enforce guidelines such as safety. Yet, alignment fails in the face of jailbreak attacks that modify inputs to induce unsafe outputs. In this paper, we present and evaluate a method to assess the robustness of LLM alignment. We observe that alignment embeds a safety classifier in the target model that is responsible for deciding between refusal and compliance. We seek to extract an approximation of this classifier, called a surrogate classifier, from the LLM. We develop an algorithm for identifying candidate classifiers from subsets of the LLM model. We evaluate the degree to which the candidate classifiers approximate the model’s embedded classifier in benign ($F_{1}$ score) and adversarial (using surrogates in a white-box attack) settings. Our evaluation shows that the best candidates achieve accurate agreement (an $F_{1}$ score above 80%) using as little as 20% of the model architecture. Further, we find attacks mounted on the surrogate models can be transferred with high accuracy. For example, a surrogate using only 50% of the Llama 2 model achieved an attack success rate (ASR) of 70%–a substantial improvement over attacking the LLM directly, where we only observed a 22% ASR. These results show that extracting surrogate classifiers is a viable (and highly effective) means for modeling (and therein addressing) the vulnerability of aligned models to jailbreaking attacks.
@@ 1 Introduction
Improvements in machine learning–specifically the introduction of the transformer architectures [44]–have led to increasingly powerful large language models (LLMs). Whether they are accessed through an API or interface (e.g., GPT [32], Gemini [39], Claude [1]) or published as open-source (e.g., Llama 2 [29], Llama 3 [30], Qwen [43], etc.), these models have become the de facto tool for tasks involving natural language. They serve as foundations for new tools, in which they are augmented with new capabilities [2, 31] or fine-tuned on a downstream task [14]. In adapting these models to tasks, LLMs undergo an alignment process in which they are further trained (refined) to satisfy safety objectives [29, 36]. For example, alignment guidelines often prevent a model from outputting responses that would be offensive, discriminatory, or harmful (i.e., unsafe). It is acknowledged that alignment can (and often does) fail in adversarial settings, resulting in a jailbreak: an unsafe input being accepted as safe. Indeed, many studies have explored the automated generation of jailbreak inputs, i.e., adversarial examples [61, 25, 50, 23].
@@
We posit a novel approach to reason about LLM alignment in the context of jailbreak attacks (see  Figure 1). In this paper, we hypothesize that alignment embeds a “safety” classifier into a target model that determines whether an input prompt is safe (compliant with alignment goals) or unsafe (refusal with alignment). Consistent with this hypothesis, previous work has shown that the representations of safety queries are well separated in aligned models [24, 59], specifically in the first decoders within the LLM [21]. This work seeks to extract an approximation of the model classifier (called a surrogate classifier) and evaluate its utility in assessing a model’s robustness to adversarial inputs. In this, we explore: (a) how a surrogate classifier can be accurately extracted, and (b) how the surrogate classifier can be used to assess the safety of or attack the aligned model.
@@
We begin by exploring methods for extracting surrogate classifiers. Here, our approach systematically identifies candidates from a subset of the model architecture. Candidates are built by (a) finding a structure, i.e., a subset of the architecture of the model, that best represents the embedded safety classifier, (b) adding a classification head, and (c) training this classification head on unsafe and safe inputs to map the features extracted from the structure to the model predictions (positive for refusal, negative for compliance). The candidates are then evaluated in benign and adversarial settings. For the benign settings, we measure how well the candidate classifier agrees with the model’s safety classifier on safe and unsafe inputs. For the adversarial settings, we measure transferability of adversarial inputs between the target model and candidate classifiers to determine how well the latter capture the decision of the embedded safety classifier.
@@
We perform a detailed analysis of the efficacy of our approach over 5 models and 2 datasets. We first evaluate candidates with respect to benign inputs, assessing whether each candidate classifier agrees with the safety classification of the LLM model. We find that the best candidates achieve accurate agreement (an $F_{1}$ score above 80%) using as little as 20% of the model architecture. Next, we evaluate the candidate classifiers with respect to transferability of adversarial inputs (i.e., as generated using best-in-class jailbreaking algorithms) from the candidates onto the LLM. These latter experiments show that some candidate classifiers are extremely effective tools for targeting LLMs. For example, a surrogate classifier using only 50% of the Llama 2 model achieved an attack success rate (ASR) of 70%–a substantial improvement over attacking the LLM directly, where we observed a 22% ASR.
@@
Our results suggest that investigating the robustness can be furthered (and in many cases made more efficient) by extracting surrogate classifiers that target alignment tasks. In this paper, we have found that approximations of LLMs may be sufficient to explore the security of their deployment–and at the same time offer potential vectors for developing more effective attacks on alignment.
@@
  • We explore the existence and structure of safety classifiers in LLMs. We show how these classifiers are used to decide whether an input complies with the alignment.
  • We develop algorithms to extract highly accurate surrogate safety classifiers from LLMs by selectively extracting and refining subsets of LLM architectures.
  • We evaluate the candidate classifiers with respect to their size and the ability to transfer attacks onto the victim LLM. We show that transferred attacks are more successful than direct attacks on the model in several settings.
@@ 2.1 Large Language Models (LLMs)
Large language model. A LLM models the conditional probability of the next token (a unit of text for the model) with respect to previous tokens. Let $V$ be the vocabulary of the tokens and $V^{\ast}=\bigsqcup\limits_{n=1}^{N}V^{n}$ the corresponding input space of sequences, where $N$ denotes the context window of the model (maximum amount of tokens that can be processed by the model). Given a sequence of tokens $x=x_{1}x_{2}\dots x_{T}\in V^{\ast}$, the model aims to produce the next token $x_{T+1}$ by approximating the probability distribution.
@@
Chat template. By construction, LLMs simply predict the next token. Thus, before deployment, the model undergoes instruction tuning: a process which encodes “roles” in the model such that it can interact with a user. Part of this process includes a chat template: a fixed modification of the input to dissociate the roles. It relies on special tokens from the model to enforce a format. As a result of different training settings, different models generally come with different chat templates. Formally, for a given model, there exists a template $(x_{\text{pre}},x_{\text{post}})\in V^{\ast}\times V^{\ast}$ such that any input text $x$ is preprocessed by prepending $x_{\text{pre}}$ and appending $x_{\text{post}}$.
@@ 2.2 Alignment and Safety
Alignment. Since the training data used for LLMs comes from diverse sources, biases and undesirable behaviors emerge [11]. In order to prevent those behaviors from happening, models often go through an alignment process that regulates their outputs according to given guidelines. Several methods exist and are not mutually exclusive [29]. Approaches such as supervised fine-tuning (SFT) uses sample answers from humans, while reinforcement learning with human feedback (RLHF) [3, 7] trains a neural network that acts as a reward for the LLM. Since training another reward model may be costly, direct preference optimization (DPO) [36] uses the LLM as its own reward model.
Guidelines and taxonomies. The existence of alignment as a technique is a direct product of the complexity of identifying what is precisely expected from models (i.e., their objective). Safety is a broad term, thus, previous work has identified taxonomies [16, 47] that represent a broad range of unsafe behaviors. This allows researchers to evaluate alignment on a finer scale and understand where it could be improved. For example, when the prompts are decomposed into different categories such as self-harm, privacy, harassment, etc., models tend to refuse the prompts related to self-harm and accept the prompt related to privacy [9].
Identifying unsafe outputs. Similarly to their inputs, classifying LLMs outputs is challenging. Since aligned models tend to create strong refusal answers for detected unsafe prompts, a naive solution is to test whether certain refusal keywords are in the output (e.g., "Sorry", "as a responsible AI", etc.) [61]. Alternatively, judge models (also LLMs) are trained to classify the safety of outputs with more [16] or less [27] granularity on the classes they predict, similarly to the taxonomies explained above. A limitation of judge models is that they can also be attacked [26, 37], leading to a possible high rate of false negatives (unsafe outputs judged as safe).
@@ 2.3 Attacks on LLMs
LLMs, like any class of machine learning models, are prone to attacks. While several exploits exist (e.g., model-stealing [4], text misclassification, etc. [58]), the jailbreak exploit remains the most prevalent. Jailbreak refers to a compliance from an aligned LLM for an unsafe query, as defined by the alignment guidelines. Jailbreak attacks fall into one of two threat models: the white-box threat model, in which model parameters are known, and the black-box threat model, in which only access to the output of the model is assumed.
White-box. These attacks have been studied on open-weights models such has Llama 2 [29], Llama 3 [30], or Vicuna [6]. The most prevalent, GCG [61], uses greedy coordinate gradient (GCG) to mitigate the absence of mapping from embeddings (i.e., vectors) to tokens. Following this, many new white-box algorithms have been introduced [48, 8], improving existing algorithms [52, 34, 23, 24], using different objectives [54], or using specific properties of the model (e.g., logits [22], special tokens [55], or generation [15]. In terms of efficacy, attacks like GCG have an average attack success rate (ASR) above 50% [27] on most open-source models.
Black-box. These attacks have been widely studied, given their prevalence in practical scenarios [32, 1]. Most of them are driven by empirical heuristics: they apply transformations to the initial prompt such as ASCII-art [18], translation into underrepresented languages [49], encoding and decoding [12], or using pre-existing templates [51]. Black-box attacks achieve remarkable efficacy. For instance, GPTFuzzer [50] can achieve ASR as high as 90% on commercial LLMs such has ChatGPT.
@@ 3 Methodology
This section outlines our approach for extracting a surrogate classifier from LLMs, which serves as the basis for our experiments. Before we detail our methodology, we first explain a preliminary experiment whose results inform our approach.
@@ 3.1 Preliminaries
As explained previously, LLMs are aligned but are prone to attacks. Regardless of the threat model, jailbreak attacks have given limited understanding of where alignment fails since they involve heuristics in either the adversarial objective or the perturbation they apply to the input. In this work, we hypothesize that alignment embeds a safety classifier in the model, whose existence we outline next. By extracting this safety classifier, we can create a new class of optimization-based attacks that reduces jailbreaking to misclassification on this surrogate classifier.
@@
Motivation of existence. Refusal from aligned models often comes with a set of predefined outputs (e.g., “I cannot fulfill your request. I’m just an AI”) as a product of alignment. As the space of refusal outputs is very limited, such outputs can be easily detected. We assume the following: the first generated token after the input is enough information to know whether the output will be a refusal or not. Under this assumption, there exists a set of refusal tokens $V_{r}$ such that, for a given input sequence of tokens $x=x_{1}x_{2}\dots x_{T}\in V^{\ast}$, $x_{T+1}\in V_{r}$ if and only if the output answer is a refusal. Thus, classifying unsafe inputs reduces to verifying $x_{T+1}\in V_{r}$. In a deterministic generation (i.e., without sampling), this is done by finding the token $x_{T+1}$ that satisfies $x_{T+1}=\arg\max\limits_{i\in V}\text{softmax}(Ah^{(D)}_{T})_{i}=\arg\max\limits_{i\in V}[Ah^{(D)}_{T}]_{i}$. As explained in Section 2.1, the last embedding (after the last decoder) $h_{T}^{(D)}$ entirely determines the next token. Moreover, the sequence of decoders in the model transforms the sequence of embeddings such that for a given decoder $i$, the $t$-th element $h_{t}^{(i)}$ is a function of $h_{1}^{(i-1)},\dots,h_{t}^{(i-1)}$. Thus, the partial information on the classification contained in $h_{1}^{(i-1)},\dots,h_{T}^{(i-1)}$ propagates to $h_{T}^{(i)}$ and so forth until it reaches $h_{T}^{(D)}$.
@@
Separation metric. For the classifier to exist, there must be separation between the two classes, i.e., inputs classified as unsafe (refusal) and inputs classified as safe (compliance). To demonstrate this, we measure the silhouette score on the embeddings associated to the two types of inputs (safe and unsafe). We use the silhouette score to determine how well clustered the two safety classes are and evaluate if there is indeed a separation, the metric is defined as follows:
$s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$ (2)
@@
where $a(i)$ denotes the mean intra-cluster distance (average distance to all other points in the same cluster) and $b(i)$ the mean nearest-cluster distance (average distance to points in the nearest cluster).
@@
This score ranges from $-1$ to $1$ where scores below 0 mean that unsafe and safe representations overlap, and scores above 0.25 or 0.5 imply a weak or reasonable separation [38].
Figure 2: Silhouette score of the embeddings for different models on two datasets.
@@
Result. We measure this score on five LLMs and two datasets (see Section 4.1 for the detailed experimental setup). Figure 2 shows the silhouette score of the embeddings for different models and datasets, with respect to the normalized decoder position (ratio between the current decoder position and the total number of decoders).
@@
While it would be expected to see a nondecreasing trend, we see that after the silhouette score reaches its maximum, it starts to decrease. A possible explanation for this phenomenon is rooted in the fact that models are designed to perform multiple tasks. Therefore, after extracting and integrating the safety information from the input, the other tasks add more information to the embedding that is not related to safety. Thus, the safety separation decreases, but remains high enough for the classification. Besides the difference in degree of separation between the two datasets, we note that the trend is exactly the same.

### v2

@@ Abstract
Alignment in large language models (LLMs) is used to enforce guidelines such as safety. Yet, alignment fails in the face of jailbreak attacks that modify inputs to induce unsafe outputs. In this paper, we introduce and evaluate a new technique for jailbreak attacks. We observe that alignment embeds a safety classifier in the LLM responsible for deciding between refusal and compliance, and seek to extract an approximation of this classifier: a surrogate classifier. To this end, we build candidate classifiers from subsets of the LLM. We first evaluate the degree to which candidate classifiers approximate the LLM’s safety classifier in benign and adversarial settings. Then, we attack the candidates and measure how well the resulting adversarial inputs transfer to the LLM. Our evaluation shows that the best candidates achieve accurate agreement (an $F_{1}$ score above 80%) using as little as 20% of the model architecture. Further, we find that attacks mounted on the surrogate classifiers can be transferred to the LLM with high success. For example, a surrogate using only 50% of the Llama 2 model achieved an attack success rate (ASR) of 70% with half the memory footprint and runtime—a substantial improvement over attacking the LLM directly, where we only observed a 22% ASR. These results show that extracting surrogate classifiers is an effective and efficient means for modeling (and therein addressing) the vulnerability of aligned models to jailbreaking attacks.
@@ 1 Introduction
Improvements in machine learning—specifically the introduction of the transformer architectures [47]—have led to increasingly powerful large language models (LLMs). Whether they are accessed through an API or interface (e.g., GPT [33], Claude [1]) or published as open-source (e.g., Llama 2 [30], Llama 3 [31], Qwen [46], etc.), these models have become the de facto tool for tasks involving natural language. They serve as foundations for new tools, in which they are augmented with new capabilities [2, 32] or fine-tuned on a downstream task [16]. In adapting these models to tasks, LLMs undergo an alignment process in which they are further trained (refined) to satisfy safety objectives [30, 39]. For example, alignment guidelines often prevent a model from outputting responses that would be offensive, discriminatory, or harmful (i.e., unsafe). It is acknowledged that alignment can (and often does) fail in adversarial settings, resulting in a jailbreak: an unsafe input being accepted as safe. Indeed, many studies have explored the automated generation of jailbreak inputs, i.e., adversarial examples [65, 27, 25]. However, these attacks, specifically in white-box settings, suffer from two major drawbacks: they are inefficient (i.e., high runtime and memory footprint) and employ various heuristics (e.g., maximize the likelihood of a specific output sentence [65], apply fixed changes to the input [53, 20], etc.) that limit the search space of adversarial inputs and thus lower their efficacy. Thus, these flaws limit the potential of attacks to assess the robustness of models.
@@
In this paper, we posit a novel approach to reason about LLM alignment in the context of jailbreak attacks (see  Figure 1). Here, we hypothesize that alignment embeds a “safety” classifier into an LLM that determines whether an input prompt is safe (compliant with alignment goals) or unsafe (refusal with alignment). In order to evaluate this hypothesis, we extract an approximation of the model’s classifier (called a surrogate classifier) and evaluate its utility in assessing a model’s robustness to adversarial inputs. We explore how a surrogate classifier can be extracted, and thereafter how the surrogate classifier can be used to assess robustness. In this, we demonstrate that surrogate classifiers are an efficient and accurate means of evaluating the robustness of a target LLM.
@@
We begin by exploring methods for extracting surrogate classifiers. Here, our approach systematically identifies candidates from subsets of the LLM’s architecture. Candidates are built by (a) finding a structure, i.e., a subset of the architecture of the model, that best represents the embedded safety classifier, (b) adding a classification head, and (c) training this classification head on unsafe and safe inputs to map the features extracted from the structure to the model predictions (positive for refusal, negative for compliance). The candidates are then evaluated in benign and adversarial settings. For the benign settings, we measure how well the candidate classifier agrees with the model’s safety classifier on safe and unsafe inputs. For the adversarial settings, we measure transferability of adversarial inputs between the LLM and candidate classifiers to determine how well the latter capture the decision of the embedded safety classifier.
@@
We perform a detailed analysis of the efficacy of our approach over 4 models and 2 datasets. We first evaluate candidates with respect to benign inputs, assessing whether each candidate classifier agrees with the safety classification of the LLM model. We find that the best candidates achieve accurate agreement (an $F_{1}$ score above 80%) using as little as 20% of the model architecture. Next, we evaluate the candidate classifiers with respect to transferability of adversarial inputs (i.e., as generated using best-in-class jailbreaking algorithms) from the candidates onto the LLM. These latter experiments show that some candidate classifiers are extremely effective and efficient tools for targeting LLMs. For example, a surrogate classifier using only 50% of the Llama 2 model achieved an attack success rate (ASR) of 70%—a substantial improvement over attacking the LLM directly, where we observed a 22% ASR—with less than half the runtime (350s/sample against 850s/sample) and the memory footprint (14 GB of VRAM usage against 31 GB).
@@
Our results suggest that investigating the robustness can be furthered (and in many cases made more efficient) by extracting surrogate classifiers that target alignment tasks. In this paper, we have found that approximations of LLMs may be sufficient to explore the security of their deployment, and at the same time offer potential vectors for developing more effective attacks on alignment.
@@
  • We introduce a novel paradigm for jailbreak attacks that extracts a surrogate classifier from a subset of the LLM’s weights and attacks it.
  • We evaluate this paradigm on several state-of-the-art LLMs, showing that attacking the surrogate classifier and transferring to the LLM is more successful than attacking the LLM directly.
  • We show that studying alignment robustness is scalable, as our approach significantly reduces the memory footprint and runtime of the attack.
@@ 2.1 Large Language Models (LLMs)
Large language model. An LLM models the conditional probability of the next token (a unit of text for the model) with respect to previous tokens. Let $V$ be the vocabulary of the tokens and $V^{\ast}=\bigsqcup\limits_{n=1}^{N}V^{n}$ the corresponding input space of sequences, where $N$ denotes the context window of the model (maximum amount of tokens that can be processed by the model). Given a sequence of tokens $x=x_{1}x_{2}\dots x_{T}\in V^{\ast}$, the model aims to produce the next token $x_{T+1}$ by approximating the probability distribution.
@@ 2.2 Alignment and Safety
Alignment. Since the training data used for LLMs comes from diverse sources, biases and undesirable behaviors emerge [14], leading to misaligned pretrained models [3]. In order to prevent those behaviors from happening, models often go through an alignment process that regulates their outputs according to given guidelines. Several methods exist and are not mutually exclusive [30]. Approaches such as supervised fine-tuning (SFT) use human-generated sample responses, while reinforcement learning with human feedback (RLHF) [4, 8] trains a neural network that acts as a reward for the LLM. Since training another reward model may be costly, direct preference optimization (DPO) [39] uses the LLM as its own reward model.
Guidelines and taxonomies. The existence of alignment as a technique is a direct product of the complexity of identifying what is precisely expected from models (i.e., their objective). Safety is a broad term; therefore, previous work has identified taxonomies [18] that represent a wide range of unsafe behaviors. This allows researchers to evaluate alignment on a finer scale and understand where it could be improved. For example, when the prompts are decomposed into different categories such as self-harm, privacy, harassment, etc., models tend to refuse the prompts related to self-harm and accept the prompt related to privacy [10].
Identifying unsafe outputs. Classifying LLM outputs is challenging. Since aligned models tend to create strong refusal responses for detected unsafe prompts, a naive solution is to test whether certain refusal keywords are in the output (e.g., ”Sorry”, ”as a responsible AI”, etc.) [65]. Alternatively, judge models (also LLMs) are trained to classify the safety of outputs with more [18] or less [29] granularity on the classes they predict, similarly to the taxonomies explained above. A limitation of judge models is that they can also be attacked [28, 40], leading to a possible high rate of false negatives (unsafe outputs judged as safe).
@@ 2.3 Attacks on LLMs
LLMs, like any machine learning model, are prone to attacks. While several exploits exist (e.g., model-stealing [5], text misclassification, etc. [61]), the jailbreak exploit remains the most prevalent. Jailbreak refers to the compliance of an aligned LLM for an unsafe query, as defined by the alignment guidelines. Jailbreak attacks fall into one of two threat models: the white-box threat model, in which model parameters are known, and the black-box threat model, in which only access to the output of the model is assumed.
White-box. These attacks have been studied on open-weights models such has Llama 2 [30], Llama 3 [31], or Vicuna [7]. The most prevalent, GCG [65], uses greedy coordinate gradient (GCG) to mitigate the absence of mapping from embeddings (i.e., vectors) to tokens. Following this, many new white-box algorithms have been introduced [52, 9], improving existing algorithms [55, 37, 25, 26], using different objectives [57], or using specific properties of the model (e.g., logits [24], special tokens [58], or generation [17]). Attacks like GCG have an average attack success rate (ASR) above 50% [29] on most open-source models.
Black-box. These attacks have been widely studied, given their prevalence in practical scenarios [33, 1]. Most of them are driven by empirical heuristics: they apply transformations to the initial prompt such as ASCII-art [20], translation into underrepresented languages [53], or using pre-existing templates [54]. Black-box attacks achieve remarkable efficacy. For instance, ArtPrompt [20] can achieve ASR as high as 80% on commercial LLMs such has GPT-3.5.
@@ 3 Methodology
This section outlines our approach for extracting and attacking surrogate classifiers of LLMs. Before we detail our methodology, we first explain a preliminary experiment whose results inform our approach.
@@ 3.1 Preliminaries
As explained previously, LLMs are aligned but are prone to attacks. Regardless of the threat model, jailbreak attacks have given limited understanding of where alignment fails, since they involve heuristics in either the adversarial objective or the perturbation they apply to the input. In this work, we hypothesize that alignment embeds a safety classifier in the model, whose existence we outline next. By extracting this safety classifier, we can create a new class of optimization-based attacks that reduces jailbreaking to misclassification on this surrogate classifier.
@@
Motivation of existence. Refusal from aligned models often comes with a set of predefined outputs (e.g., “I cannot fulfill your request. I’m just an AI”) as a product of alignment. As the space of refusal outputs is very limited, such outputs can be easily detected. We assume the following: the first generated token after the input is enough information to know whether the output will be a refusal or not. Under this assumption, there exists a set of refusal tokens $V_{r}$ such that, for a given input sequence of tokens $x=x_{1}x_{2}\dots x_{T}\in V^{\ast}$, $x_{T+1}\in V_{r}$ if and only if the output answer is a refusal. Thus, classifying unsafe inputs reduces to verifying $x_{T+1}\in V_{r}$. In a deterministic generation (i.e., without sampling), this is done by finding the token $x_{T+1}$ that satisfies $x_{T+1}=\arg\max\limits_{i\in V}\text{softmax}(Ah^{(D)}_{T})_{i}=\arg\max\limits_{i\in V}[Ah^{(D)}_{T}]_{i}$. As explained in Section 2.1, the last embedding (after the last decoder) $h_{T}^{(D)}$ entirely determines the next token. Moreover, the sequence of decoders in the model transforms the sequence of embeddings such that for a given decoder $i$, the $t$-th element $h_{t}^{(i)}$ is a function of $h_{1}^{(i-1)},\dots,h_{t}^{(i-1)}$. Thus, the partial information on the classification contained in $h_{1}^{(i-1)},\dots,h_{T}^{(i-1)}$ propagates to $h_{T}^{(i)}$ until it reaches $h_{T}^{(D)}$.
@@
Separation metric. For the classifier to exist, there must be separation between the two classes, i.e., inputs classified as unsafe (refusal) and inputs classified as safe (compliance). To demonstrate this, we measure the average silhouette score on the embeddings associated to the safe and unsafe inputs. We use the silhouette score to determine how well the two safety classes cluster and to evaluate whether there is a separation, defined as follows for a given data point $i$:
$s(i)=\frac{d_{\text{inter}}(i)-d_{\text{intra}}(i)}{\max(d_{\text{intra}}(i),d_{\text{inter}}(i))}$ (2)
@@
where $d_{\text{intra}}(i)$ denotes the mean intra-cluster distance (average distance to all other points in the same cluster) and $d_{\text{inter}}(i)$ the mean nearest-cluster distance (average distance to points in the nearest cluster).
@@
This score ranges from 1-1 to $1$ where scores below 0 indicate that representations overlap, and scores above 0.25 or 0.5 imply a weak or reasonable separation [41].
Figure 2: Silhouette score (measure of separation) of unsafe and safe input embeddings for different LLMs.
@@
Result. We measure this score on four LLMs and two datasets (see Section 4.1 for the detailed experimental setup). Figure 2 shows the silhouette score of the embeddings for different models and datasets, with respect to the normalized decoder position (ratio between the current decoder position and the total number of decoders).
@@
While it would be expected to see a nondecreasing trend [^1] Usually, the subsequent layers of classifiers do not lose information on the classification, hence a nondecreasing trend. For LLMs, this classification is only one of many tasks that are performed, which is why we observe a compression of the linear information in later layers., we see that after the silhouette score reaches its maximum, it starts to decrease. A possible explanation for this phenomenon is rooted in the fact that models are designed to perform multiple tasks. Therefore, after extracting and integrating the safety information from the input, the other tasks add more information to the embedding that is not related to safety. Thus, the safety separation decreases (i.e., the linear information is compressed), but remains high enough for the classification. Besides the difference in magnitude of separation between the two datasets, we note that the trend is exactly the same for all models.